Five years ago, Tom was one third as old as his father was then. In 5 years, Tom will be half as old as his father will be then. Find their ages now.
Answer by scott8148(6628)

let t="Tom now", and f="father now"

"Five years ago, Tom was one third as old as his father was then"
... 3(t-5) = f-5
distributing ... 3t-15 = f-5
adding 5 ... 3t-10 = f

"In 5 years, Tom will be half as old as his father will be then"
... 2(t+5) = f+5
distributing ... 2t+10 = f+5
subtracting 5 ... 2t+5 = f

substituting ... 3t-10 = 2t+5
adding 10-2t ... t = 15

substituting ... 2(15)+5 = f ... 35 = f